A chemist is studying the following equilibirum, which has the given equilibrium constant at a certain temperature:

3O2 -- 2O3

Kp= 1.0 x10. ^-7

He fills a reaction vessel at this temperature with 11.atm of oxygen gas. Use this data to answer the questions in the table below.

-Can you predict the equilibrium pressure of O3, using only the tools available to you within ALEKS?

-If you said yes, then enter the equilibrium pressure of O3 at right. Round your answer to 1 significant digit.
Answer #1

PO2 = 11 atm

Kp = 1.0 x 10-7

The equilibrium reaction is as follows:

3O2 2O3

3O2 2O3
I(atm) 11 0
C(atm) -3x +2x
E(atm) 11-3x 2x

The Kp expression is as follows:

Kp = [O3]2 / [O2]3

As per the ICE chart.

1.0 x 10-7 = [2x]2 / [11-3x]3

1.0 x 10-7 = [2x]2 / [11]3 --------- [11-3x] 11, since x<<< 11

1.0 x 10-7 = 4x2 / [1331]

x = 0.005768

As per the ICE chart,

[O3]eq = 2x = 2 x 0.005768

[O3]eq = 0.01154

Thus,

equilibrium pressure of O3 is 0.01 atm [1 S.F]